# Provable Scaling Laws of Feature Emergence from Learning Dynamics of Grokking

### v4

@@ 5 Stage III: Interactive feature learning
The starting point of Stage II is to simplify the exact backpropagated gradient $G_{F}=P_{\eta}\tilde{Y}\tilde{Y}^{\top}\tilde{F}B$ (Eqn. 5) with $B:=(\tilde{F}^{\top}\tilde{F}+\eta I)^{-1}$ to $G_{F}\propto\eta\tilde{Y}\tilde{Y}^{\top}F$, by two approximations: (1) $B\propto I$, and (2) $P_{\eta}\propto\eta I$. The two approximations are valid due to Thm. 1 when the hidden weights $W$ is randomly initialized. When training continues, $W$ evolves from random initialization and the conditions may not hold anymore. In this section we put them back and study their behaviors.
@@ 5.3 Diversity enhancement with Muon
Note that Eqn. 11 is closely related to $\mathcal{E}$, under the assumption of homogeneous/reversible activation, i.e., $\sigma(x)=C\sigma^{\prime}(x)x$ with a constant $C$ (Zhao et al., 2024; Tian et al., 2020). In such setting, Eqn. 7 is related to the gradient dynamics with a PSD matrix $A(\mathbf{w})=X^{\top}D(\mathbf{w})\tilde{Y}\tilde{Y}^{\top}D(\mathbf{w})X$.
@@ 6 Extension to deeper architectures
Stage I. Stage I does not change since $F_{L}$ is still a random representation. Then $V$ starts to learn and converges to ridge solution (Eqn. 4), the backpropagated gradient $G_{L}$ now carries meaningful information: $G_{L}\propto\tilde{Y}\tilde{Y}^{\top}F_{L}$ (Eqn. 6), which initiates Stage II.
@@ 7 Discussion
Several ways exist to ensure this nonzero gradient. Regularization mechanisms, such as weight decay, help guarantee the presence of a nonzero $G_{F}$, and conveniently allow us to demonstrate the feature learning process in a mathematically rigorous way. However, we want to emphasize that regularization is not the only option. For example, $G_{F}$ might become large during lazy learning and later vanish due to perfect overfitting (see Eqn. 5). Notably, (Prieto et al., 2024) replaces the standard softmax (exponential term) with a stable softmax (linear term), which from the perspective of the Li2 framework, slows down overfitting to the labels. This maintains a nonzero backpropagated gradient, thereby preserving the potential for subsequent feature learning.
@@ 9 Conclusion, Limitations, and Future Work
Optimizing in Real domain. The above analysis uses complex irreducible representations. For real $\mathbf{w}$, $\hat{S}_{k,r}$ will be a complex conjugate of $\hat{S}_{-k,r}$ for conjugate irreducible representations $k$ and k-k. This means that we can partition the sum in Eqn. 21 into real and complex parts:
@@
For Abelian group $H$ with $|H|=M>2$, all irreducible representations are 1-dimensional, and at least one of it is complex. Since $\mathbb{C}$ is treated as 2D space in optimization, it has at least 1 degree of freedom to change without changing its function value (Eqn. 26). So the Hessian has at least 1 zero eigenvalue. For non-Abelian group, there is at least one irreducible representation $k$ with dimension greater than 1, which means it has at least 1 degrees of freedom to change $\hat{S}_{k,r}$ without changing $|\sum_{r}\operatorname{tr}(\hat{S}_{k,r})|^{2}$ and thus its function value (Eqn. 26). So the Hessian has at least 1 zero eigenvalue. ∎
@@
Stability of local maxima. For the quadratic case (after mean removal), with the collinear and equal length $\mathbf{u}$ and $\mathbf{v}$ required by local maxima, $\mathcal{E}$ can be written as a positive semidefinite quadratic in the block masses $c_{k}$ (Eqn. 23):
@@ A.4 Memorization
where $\mathbf{z}=[z_{h}]=[\langle R_{h},S\rangle_{F}]\in\mathbb{C}^{M}$ defined in Eqn. 14. Computing each row $k$ in $\Phi_{\mathcal{S}}^{\ast}\mathbf{z}$ and use the property of projection matrix $\Pi_{k}$ (Eqn. 36), we have:
@@
which is exactly the same form as the decomposition (Eqn. 21) in Thm. 2 (but a much cleaner derivation). Therefore, all the local maxima of $\mathcal{E}_{\mathcal{S}}$ are still in the same form as Thm. 2, but we just remove those local maxima that are in isotype/irreps $k\in\mathcal{S}$, and focus on missing ones. ∎
@@
Lemma 8 (The properties of the dynamics in Eqn. 11).

### v5

@@ 5 Stage III: Interactive feature learning
The starting point of Stage II is to simplify the exact backpropagated gradient $G_{F}=P_{\eta}\tilde{Y}\tilde{Y}^{\top}\tilde{F}B$ (Eqn. 4) with $B:=(\tilde{F}^{\top}\tilde{F}+\eta I)^{-1}$ to $G_{F}\propto\eta\tilde{Y}\tilde{Y}^{\top}F$, by two approximations: (1) $B\propto I$, and (2) $P_{\eta}\propto\eta I$. The two approximations are valid due to Thm. 1 when the hidden weights $W$ is randomly initialized. When training continues, $W$ evolves from random initialization and the conditions may not hold anymore. In this section we put them back and study their behaviors.
@@ 5.3 Diversity enhancement with Muon
Note that Eqn. 10 is closely related to $\mathcal{E}$, under the assumption of homogeneous/reversible activation, i.e., $\sigma(x)=C\sigma^{\prime}(x)x$ with a constant $C$ (Zhao et al., 2024; Tian et al., 2020). In such setting, Eqn. 6 is related to the gradient dynamics with a PSD matrix $A(\mathbf{w})=X^{\top}D(\mathbf{w})\tilde{Y}\tilde{Y}^{\top}D(\mathbf{w})X$.
@@ 6 Extension to deeper architectures
Stage I. Stage I does not change since $F_{L}$ is still a random representation. Then when $V$ starts to learn and converges, the backpropagated gradient $G_{L}$ now carries meaningful information: $G_{L}\propto\tilde{Y}\tilde{Y}^{\top}F_{L}$ (Eqn. 5), which initiates Stage II.
@@
7 When does grokking happen?
We first want to clarify that grokking (i.e., delayed generalization) is a phenomenon/observation that may suggest multiple underlying situations. When we observe “grokking happens”, it is often the case that the training gets stuck in lazy learning for a while before moving to feature learning. On the other hand, when we observe ”grokking does not happen”, then we may fall into one of the three very different situations: (1) training simply gets stuck in lazy training, (2) feature learning happens very early and the gap between training and test performance is always small, or (3) feature learning has converged to non-generalizable solutions due to e.g., insufficient or noisy data, and thus test performance is never improved. Note that (2) is a very good outcome while (1) and (3) are bad ones.
Therefore, it would be tricky to directly relate the hyperparameters with grokking behaviors. Instead, using Li2 framework, we focus on relating them with the backpropagated gradient $G_{F}$ and the underlying feature learning process, which is more fundamental. Only when the $G_{F}$ become large and sends a sufficient “amount” of the key term $\tilde{Y}\tilde{Y}^{\top}F$ down to the hidden layers, grokking could happen. Here we categorize these factors into several categories.
Learning rate.  (Gromov, 2023) reports that grokking happens without regularization, but with a large initial learning rate (verified by the author). This corresponds to increasing the strength of $G_{F}(t)\propto t\tilde{Y}\tilde{Y}^{\top}F$ at the initial phase of learning so that the hidden layers receives enough correct gradient signal to start the feature learning.
Stay longer before overfitting.  (Prieto et al., 2025) uses stable softmax (linear form) rather than regular softmax (exponential form) in computing probability. This prevents the model from overfitting to the label too quickly, and thus maintains a nonzero backpropagated gradient that can be useful for feature learning. (Kumar et al., 2024) also reports that grokking happens without regularization, using vanilla SGD optimizer. Our explanation is that it may take longer for SGD to converge to $V_{\textrm{ridge}}$ than Adam, and during that period, the hidden layer has already accumulated a sufficient amount of correct gradient signal.
Weight initialization.  (Liu et al., 2023) reports that grokking happens with small initialization, regardless of the weight decay. This is straightforward from our framework, since $G_{F}(t)=O(\alpha)+t\tilde{Y}\tilde{Y}^{\top}F+O(\alpha t)+O(t^{2})$ and if the weight initialization $\alpha$ is small, then $G_{F}(t)$ is dominated by clear signal term $t\tilde{Y}\tilde{Y}^{\top}\tilde{F}$, which leads to fast feature learning. If $\alpha$ is large, then $O(\alpha)$ term is large and the initial phase of $G_{F}$ contains too much noise, and we need to rely on the signal provided by the convergence phrase of $G_{F}$ controlled by the weight decay $\eta$. This is consistent with the finding by (Liu et al., 2023) that for large weight initialization, regularization is needed for grokking to happen, and small regularization leads to slow grokking transition.
Scaling factor $\beta$ of the output.  (Kumar et al., 2024; Chizat et al., 2019) reports that scaling the output by a factor $\beta>1$ will make the grokking faster. From Li2 framework, this corresponds to optimizing $J_{\beta}(V)=\|\tilde{Y}-\beta\tilde{F}V\|^{2}_{F}+\eta\|V\|^{2}_{F}$. Following a similar derivation as in Sec. C, in Sec. C.3 we can show that at the initial phase, the backpropagated gradient $G_{F}(t)=O(\alpha)+t\left[\beta\tilde{Y}\tilde{Y}^{\top}\tilde{F}+O(\beta^{2}\alpha)+O(\beta^{3}\alpha^{2})\right]+O(t^{2})$. So if $\beta>1$ is large and the initialization $\alpha\ll 1$ then the signal term $t\beta\tilde{Y}\tilde{Y}^{\top}\tilde{F}$ becomes more dominant than the case of $\beta=1$, and the grokking happens faster. However, if we have too large $\beta$ over $\alpha$, then the term $O(\beta^{2}\alpha)$ may cause trouble in feature learning and the final $G_{F}(+\infty)$ becomes small (Eqn. 69). Equivalently, same theory applies if we downscale the target label (Kumar et al., 2024).
Weight decay $\eta$. According to Eqn. 4, since $G_{F}(+\infty)\propto\eta\tilde{Y}\tilde{Y}^{\top}F$, it is clear that the weight decay $\eta$ becomes the learning rate of feature learning process. This coincides with findings in empirical works (Power et al., 2022; Clauw et al., 2024) that low regularization leads to slow grokking transition. This is also consistent with $t\sim 1/\eta$ laws to start grokking (Liu et al., 2023) or reach maximal test performance (Lewkowycz & Gur-Ari, 2020).
Data size $n$. Our sample analysis (Theorem. 4) shows the local maxima can be kept with sufficient number of samples ($n\gtrsim M\log M$). Intuitively, more samples lead to better shaped local maxima with less noise and thus the feature learning is faster, which is consistent with (Power et al., 2022; Nanda et al., 2023; Wang et al., 2024a; Gromov, 2023; Abramov et al., 2025).
The number of hidden nodes $K$. Our analysis requires that we need a decent number of hidden nodes $K$ to cover the diverse set of the local maxima of $\mathcal{E}$. On the other hand, Lemma 1 tells that very large $K$ may reduce $|G_{F}(+\infty)|$ and makes grokking slower. This is consistent with the finding by (Chizat et al., 2019) and NTK regime (Du et al., 2018; Jacot et al., 2018) that feature learning does not happen.
Early stopping. Model may overfit after generalization Montanari & Urbani (2025). Our framework shows that if data are scarce, generalization solutions have lower energy than memorization ones. If trained long enough, weights will move to solutions with larger $\mathcal{E}$.
@@ 8 Discussion
Several ways exist to ensure this nonzero gradient. Regularization mechanisms, such as weight decay, help guarantee the presence of a nonzero $G_{F}$, and conveniently allow us to demonstrate the feature learning process in a mathematically rigorous way. However, we want to emphasize that regularization is not the only option. For example, $G_{F}$ might become large during lazy learning and later vanish due to perfect overfitting (see Eqn. 4). Notably, (Prieto et al., 2024) replaces the standard softmax (exponential term) with a stable softmax (linear term), which from the perspective of the Li2 framework, slows down overfitting to the labels. This maintains a nonzero backpropagated gradient, thereby preserving the potential for subsequent feature learning.
@@ 10 Conclusion, Limitations, and Future Work
Optimizing in Real domain. The above analysis uses complex irreducible representations. For real $\mathbf{w}$, $\hat{S}_{k,r}$ will be a complex conjugate of $\hat{S}_{-k,r}$ for conjugate irreducible representations $k$ and k-k. This means that we can partition the sum in Eqn. 20 into real and complex parts:
@@
For Abelian group $H$ with $|H|=M>2$, all irreducible representations are 1-dimensional, and at least one of it is complex. Since $\mathbb{C}$ is treated as 2D space in optimization, it has at least 1 degree of freedom to change without changing its function value (Eqn. 25). So the Hessian has at least 1 zero eigenvalue. For non-Abelian group, there is at least one irreducible representation $k$ with dimension greater than 1, which means it has at least 1 degrees of freedom to change $\hat{S}_{k,r}$ without changing $|\sum_{r}\operatorname{tr}(\hat{S}_{k,r})|^{2}$ and thus its function value (Eqn. 25). So the Hessian has at least 1 zero eigenvalue. ∎
@@
Stability of local maxima. For the quadratic case (after mean removal), with the collinear and equal length $\mathbf{u}$ and $\mathbf{v}$ required by local maxima, $\mathcal{E}$ can be written as a positive semidefinite quadratic in the block masses $c_{k}$ (Eqn. 22):
@@ A.4 Memorization
where $\mathbf{z}=[z_{h}]=[\langle R_{h},S\rangle_{F}]\in\mathbb{C}^{M}$ defined in Eqn. 13. Computing each row $k$ in $\Phi_{\mathcal{S}}^{\ast}\mathbf{z}$ and use the property of projection matrix $\Pi_{k}$ (Eqn. 35), we have:
@@
which is exactly the same form as the decomposition (Eqn. 20) in Thm. 2 (but a much cleaner derivation). Therefore, all the local maxima of $\mathcal{E}_{\mathcal{S}}$ are still in the same form as Thm. 2, but we just remove those local maxima that are in isotype/irreps $k\in\mathcal{S}$, and focus on missing ones. ∎
@@
Lemma 8 (The properties of the dynamics in Eqn. 10).